OpenStudy (anonymous):

Is the sequence arithmetic? If so, identify the common difference. 14, 21, 42, 77, ... yes; 7 yes; –7 yes; 14 no 6.What is the 12th term of the sequence 10, 6, 2, …? –16 –24 –34 –38

OpenStudy (anonymous):

first one D.

OpenStudy (anonymous):

second C

OpenStudy (anonymous):

Am I right?

OpenStudy (whpalmer4):

Yes, both correct. First one doesn't have a common difference. Second one looks like\[a_n = 14-4n, \implies a_{12}=14-4(12) = -34\]